OpenStudy (kainui):

A fun problem, I don't know if I've asked it before here or not. Find the function such that tangent of this function is the slope of a tangent line to this function.

ganeshie8 (ganeshie8):

Deja vu!

OpenStudy (kainui):

Haha well I guess that's a yes. =P

ganeshie8 (ganeshie8):

I'll drop a hint : homonyms

OpenStudy (anonymous):

I'm not sure what you mean by "tangent of this function"?

OpenStudy (kainui):

Tangent of a function is the same as sine of the function divided by cosine of the function =P

OpenStudy (anonymous):

Hmmm interesting indeed. What kind of class is this for?

OpenStudy (kainui):

This isn't a class, just a fun question I came up one time when I was bored. It's a fairly simple differential equation that's sort of a play on words.

OpenStudy (anonymous):

Ahhh that's why I don't get it.... Haven't taken differential equations yet.

OpenStudy (kainui):

It's basically just a calculus question though, nothing too complicated. \[\Large \dfrac{dy}{dx}=\tan(y)\] Solve for y. So this is a separable equation. \[\Large \int\limits \ \frac{1}{tan(y)} dy = \int\limits dx\] since tangent is just sine over cosine, we have it's own derivative there. \[\Large \int\limits \frac{\cos (y)}{\sin (y)}dy=\int\limits dx\] Just solve these integrals and we get: \[\Large \ln| \sin(y)|=x+C\] Now we just solve for y \[\Large y=\sin^{-1}(Ke^x)\] Here I just replaced C with another arbitrary constant K=e^C. So that's all there is going on there, basic calculus 2 sort of stuff.

OpenStudy (kainui):

A common question to solve on your own is, what function is its own derivative? \[\Large y= \frac{dy}{dx}\] Now just multiply dx onto the other side and divide y to the other side to get two integrals. And then you can solve away for y. You should get something fairly predictable. =) I can explain anything you'd like to learn more of, but from what I gather you've taken at least calculus 1 correct?

OpenStudy (anonymous):

haha that's exactly correct. I'm going to be taking calc 2 next semester. I most definitely appreciate your offer of help. I will keep you in mind, friend! I can always use more help.

OpenStudy (kainui):

Yeah definitely! I like introducing people to this, so just out of curiosity, do you know what the derivative of this is? \[\Large A(r)=\pi r^2\]

OpenStudy (anonymous):

\[dA/dr=2\pi r\]Correct?

OpenStudy (kainui):

Yeah, why do you think the derivative of area is circumference?

OpenStudy (anonymous):

Because as the radius grows or shrinks, the circumference grows or shrinks in relation to the radius? Not exactly a rigorous answer lol

OpenStudy (kainui):

Well, it's kind of like an infinitesimally thin layer of area around the entire circle at that radius is what a circumference is.

OpenStudy (anonymous):

Very interesting. So as the circumference increases by dr, we add on an infinitely small ring around the circle. And this ring is nothing but a rectangle with a width of dr, and length 2 pi r. I think I understand, but not quite a solid grasp yet.

OpenStudy (anonymous):

Reminds me of solving for volume of solids of revolutions.

OpenStudy (kainui):

Yeah, exactly the same idea. So then it follows that since it has a width of dr then we can add up an infinite number of these infinitely thin slices to get Area. So you might want to try deriving the volume of a sphere, and from that derive the formula for surface area of a sphere. If that's not hard enough or you're having fun try finding the surface area and volume of a doughnut shape.

OpenStudy (anonymous):

Oh yes I have done the volume of the sphere and the doughnut shape. I found it very satisfying that the volume of the sphere came out to 4/3 pi r^3. I definitely remember now how we add on layers, as in slices or discs/washers, to the solid to find the volume. We ended the semester doing problems involving work like pumping water and with Hooke's law. I'm actually excited to start calc 2 and learn about inverse function stuff and derivatives of logarithms.

OpenStudy (kainui):

Ahh awesome. Well I'm sure you'll be fine.

OpenStudy (anonymous):

Thanks! You'll be hearing from me when my brain starts hurting.

OpenStudy (kainui):

I really like power series, integration by parts and definitely trigonometric substitution. A lot of fun stuff coming up! =)

OpenStudy (anonymous):

I hear from former students of calculus that they HATED Fourier series and the like. What do you make of that?